# How do Pre-Trained Models Support Software Engineering? An Empirical Study in Hugging Face

### v1

@@ Abstract
Open-Source Pre-Trained Models (PTMs) provide extensive resources for various Machine Learning (ML) tasks, yet these resources lack a classification tailored to Software Engineering (SE) needs. To address this gap, we derive a taxonomy encompassing 147 SE tasks and apply an SE-oriented classification to PTMs in a popular open-source ML repository, Hugging Face (HF). Our repository mining study began with a systematically gathered database of PTMs from the HF API, considering their model card descriptions and metadata, and the abstract of the associated arXiv papers. We confirmed SE relevance through multiple filtering steps: detecting outliers, identifying near-identical PTMs, and the use of Gemini 2.0 Flash, which was validated with five pilot studies involving three human annotators. This approach uncovered 2,205 SE PTMs. We find that code generation is the most common SE task among PTMs, primarily focusing on software implementation, while requirements engineering and software design activities receive limited attention. In terms of ML tasks, text generation dominates within SE PTMs. Notably, the number of SE PTMs has increased markedly since 2023 Q2. Our classification provides a solid foundation for future automated SE scenarios, such as the sampling and selection of suitable PTMs.
Index Terms:
Software Engineering Pre-trained Models, Hugging Face, Model classification
I Introduction
@@
The exponential growth of open-source platforms such as Hugging Face (HF) [1] has enhanced access to Machine Learning (ML) assets [2]. In particular, HF is distinguished by its extensive and fast-growing collection of Pre-Trained Models (PTMs)[3, 4], compared to other platforms [5][6]. In 2024 alone, HF saw an average of 2,199 new PTMs published per day and over six million daily downloads, based on our analysis of data retrieved through the platform’s public API. Moreover, it is estimated that a new repository is created every 15 seconds [7], reflecting trends observed in prior studies that reported rapid increases in repository creation and model contributions [8]. Despite this growth, the current categorization of HF resources remains largely general-purpose, often targeting ML domains as broad as computer vision or natural language processing. This generic taxonomy fails to address the specific and diverse needs of the Software Engineering (SE) community, whose requirements differ significantly from those of other fields [9]. Importantly, SE goes beyond mere coding, encompassing a set of activities structured around the well-known Software Development Life Cycle (SDLC) [10, 11]. This cycle comprises five main activities: requirements engineering, software design, software implementation, software quality assurance, and software maintenance, thereby covering the entire process from initial specification to long-term system evolution.
Other platforms like Papers with Code [12] also provide structured connections between research papers and code implementations, offering valuable categorizations of ML tasks. However, these are primarily oriented toward traditional ML domains. Even though the platform includes an SE-related category, it combines general ML tasks (e.g., reinforcement learning, semantic segmentation) with SE-specific tasks (e.g., code generation, program repair), limiting its utility for SE-focused research and practice.
@@
The absence of an SE-specific categorization in open-source model repositories [13] presents a significant barrier. Not only limits the efficient application of ML in SE, but it also hinders the continuous integration of new PTMs into a structured framework. This gap has several consequences for Automated Software Engineering (ASE) practitioners and researchers. In the absence of tailored filtering mechanisms, practitioners must manually sift through thousands of PTMs to sample those relevant for their tasks, an effort that is both time-consuming and error-prone, ultimately slowing progress and reducing reproducibility [14]. Moreover, without a task-aligned taxonomy, it becomes difficult to establish consistent benchmarks, which are essential for the advancement of automated development pipelines. More broadly, this lack of structured cataloguing prevents the SE community from systematically identifying relevant resources, evaluating progress, and enabling full automation in the multi-agent environment.
@@
Given the dynamic and evolving nature of both ML and SE [15], there is a clear need for a flexible and domain-specific classification system. Such a system would support the effective sampling, selection, and integration of PTMs tailored to the full SDLC. Prior studies have taken important steps toward organizing ML resources for SE, but each presents limitations that leave a significant gap unaddressed. Hou et al. [16] proposed a taxonomy by categorizing LLM-related papers into SE tasks and SDLC activities through a systematic literature review. However, their approach is grounded in static academic sources and lacks connection to actual open-source repositories like HF. Di Sipio et al. [13] extracted SE-relevant information from HF model cards using a semi-automated method. Yet, their validation was limited to only three PTMs, and their taxonomy does not align with the SDLC, reducing its applicability. Similarly, Yang et al. [17] curated code-focused LLMs from HF, providing useful insights into that specific subset, but their scope excludes broader SE tasks and lacks an SDLC perspective. Wang et al. [18] surveyed benchmarks for code LLMs and agents around the SDLC. However, their work does not address the classification of other types of PTMs.
@@
In this work, we address these limitations by introducing an automated classification framework for SE tasks across the SDLC. Our approach systematically maps PTMs from HF to a taxonomy of 147 SE tasks, built on prior research and established guidelines. We analyze multiple sources of descriptive content, including model card documentation, metadata, and abstracts from associated arXiv papers. The classification pipeline begins with the extraction of candidate task tokens from textual descriptions and applies several refinement steps to improve precision. These include outlier detection to remove inconsistent entries, similarity-based deduplication to eliminate near-identical variants, and an LLM support to ensure that the assigned activity reflects the model’s described purpose.
Our main contributions are as follows.
@@
    An automated classification framework that accesses the HF API, processes, filters, and classifies PTMs according to the full SDLC.
@@
    An extensive analysis of SE PTMs, covering their evolution, task distribution, model reuse, and reported attributes.
@@
    A validated, publicly available dataset of 2,205 SE PTMs.
@@
Our results offer several insights of particular relevance to the ASE community. While the number of SE PTMs in HF has grown significantly in the past few years, we find that they are heavily concentrated on the mid to late stages of the SDLC. By providing a reproducible, SDLC-aligned classification of PTMs, our work supports ASE researchers and practitioners by automating the identification and selection of the most appropriate PTMs for specific tasks. This automation reduces manual effort, improves consistency in benchmarking, and accelerates the adoption of ML-driven tools within ASE workflows, ultimately fostering more effective and targeted automation across the SDLC
@@
This article is structured as follows. Section II reviews related work on PTMs and their application in SE; Section III describes the methodology used to classify PTMs hosted on HF; Section IV presents the main results and analysis; Section V discusses the findings and explores implications; Section VI identifies threats to validity; and Section VII concludes the article and suggests future research directions.
TABLE I: Summary of relevant aspects in related work.
Study Approach [19] Data Source SDLC Auto Classification Information Used Initial Resources Identified Resources
[16] Systematic Review Published Literature 88 SE tasks across 6 SE activities Paper content 218,765 papers 395 papers
[13] Repository Mining HF dump 6 macro-tasks Model card 381,240 PTMs Tested on BERT, RoBERTa, and T5
[17] Repository Mining HF API 9 categories Model documentation 1,566 PTMs 366 PTMs
[18] Systematic Review Published Literature 30 SE tasks across 6 SE activities Academic publications 1,347 papers 181 benchmarks from 461 relevant papers
This study Repository Mining HF API + arXiv 147 SE tasks across 5 SE activities Model card description, metadata, and arXiv abstract 1,533,973 PTMs 2,205 PTMs
  • SDLC indicates alignment with the Software Development Life Cycle; Auto indicates whether the work is fully automated.
II Related Work
@@
Key aspects of the related work are summarized in Table I. A systematic literature review conducted by Hou et al. [16] analyzed 395 research papers from January 2017 to January 2024 and categorized Large Language Models (LLMs) into SE tasks. These tasks were grouped into SE activities according to the SDLC. The study identified 88 distinct SE tasks, revealing that LLMs are most commonly applied in the software development phase, with code generation and program repair being the most prevalent tasks for employing LLMs in software development and maintenance activities. In contrast, the earlier SDLC phases, such as requirements engineering and software design, are the least explored in the context of LLM applications.
@@
Di Sipio et al. [13] highlighted the lack of an SE classification of PTMs in HF, as the existing one is specific for ML. To address this gap, they defined six macro-tasks and proposed a semi-automated approach to extract SE-relevant information from model cards. However, their validation was limited to just three PTMs (BERT, RoBERTa, and T5), and their categorization did not align with the SDLC.
@@
Yang et al. [17] analyzed the ecosystem of LLMs as of August 2023, curating 366 PTMs from HF for SE. The study also explores the use of LLMs to assist in constructing and analyzing the ecosystem, which increased the model size by 16.5. They focus on code-based LLMs, which represent a more specific scope compared to LLMs for broader SE tasks.
@@
A recent study on benchmarks for code LLMs and agents [18] highlights an imbalance in their distribution across the SDLC. Around 60% of existing benchmarks focus on the software implementation phase, whereas requirements engineering and software design phases are underrepresented.
@@
Despite the notable advances mentioned above, a comprehensive SE-centric classification framework of PTMs in HF remains absent. Existing efforts, such as those by Hou et al. [16] and Wang et al. [18], are grounded in static academic literature and are not designed to analyze or update classifications based on the contents of repositories. Other repository-based studies [13, 17] do not align their categorization with the SDLC or focus narrowly on a subset of code-centric LLMs. Di Sipio et al. [13] take a step toward automation with a semi-automated mapping based on model cards, but their classification remains limited in scope and adaptability. In contrast, this paper adopts an SE perspective to extend the existing taxonomy in Hou et al. [16], to encompass a broader and more diverse set of PTMs, cataloguing them according to specific SE activities and tasks within the SDLC. By analyzing a substantially larger and broader set of PTMs, this work addresses unique SE-specific requirements not met by previous studies. As a result, we propose a novel and exhaustive classification framework tailored to the community.
III Methodology
III-A Research Objectives
@@
Following the Goal Question Metric (GQM) template [20], our goal is to analyze PTMs for the purpose of their classification with respect to their application to SE activities and tasks from the point of view of software engineers in the context of HF.
@@
This goal is structured around two RQs. The first addresses key challenges in automatic classification of PTMs, as this information is essential for the subsequent RQ:
  • RQ1: What challenges arise when automatically classifying SE PTMs in HF?
    - RQ1.1: What is the current documentation status of HF SE PTMs?
    - RQ1.2: What are the key findings regarding the correct classification of SE PTMs?
    - RQ1.3: How stable is this classification information over time?
@@
Next, we examine the role of PTMs in SE, focusing on their applications across different SE activities and tasks:
  • RQ2: How are PTMs utilized for automating SE tasks?
@@
    - RQ2.2: How are ML tasks related to SE activities?
    - RQ2.3: How are cross-activity PTMs applied across multiple SE activities?
    - RQ2.4: Which base models have been most frequently utilized for SE over time?
    - RQ2.5: What attributes characterize SE PTM in HF?
III-B Research Method
Figure 1: Classification pipeline of PTMs relevant to SE in HF.
We conducted a repository mining study [19] following the pipeline illustrated in Figure 1 to classify PTMs hosted on HF. The subsequent paragraphs detail each step of this process. This approach supports reproducibility, allowing anyone with access to the replication package [21] to validate the results. Additionally, the pipeline is designed to be easily updated, so researchers can reapply it to classify new PTMs added to the platform using the same method. The inclusion criteria are summarized in Table II.
TABLE II: Inclusion criteria to identify SE PTMs in HF.

### v2

@@ Abstract
Context: Open-source Pre-Trained Models (PTMs) provide extensive resources for various Machine Learning (ML) tasks, yet these resources lack a classification tailored to Software Engineering (SE) needs to support the reliable identification and reuse of models for SE. Objective: To address this gap, we derive a taxonomy encompassing 147 SE tasks and apply an SE-oriented classification to PTMs in a popular open-source ML repository, Hugging Face (HF). Method: Our repository mining study followed a five-phase pipeline: (i) identification SE tasks from the literature; (ii) collection of PTM data from the HF API, including model card descriptions and metadata, and the abstracts of the associated arXiv papers; (iii) text processing to ensure consistency; (iv) a two-phase validation of SE relevance, involving humans and LLM assistance, supported by five pilot studies with human annotators and a generalization test; (v) and data analysis. This process yielded a curated catalogue of 2,205 SE PTMs. Results: We find that most SE PTMs target code generation and coding, emphasizing implementation over early or late development stages. In terms of ML tasks, text generation dominates within SE PTMs. Notably, the number of SE PTMs has increased markedly since 2023 Q2, while evaluation remains limited: only 9.6% report benchmark results, mostly scoring below 50%. Conclusions: Our catalogue reveals documentation and transparency gaps, highlights imbalances across SDLC phases, and provides a foundation for automated SE scenarios, such as the sampling and selection of suitable PTMs.
1 Introduction
@@
The exponential growth of open-source platforms such as Hugging Face (HF) huggingface has enhanced access to Machine Learning (ML) assets zhao2024empirical. In particular, HF is distinguished by its extensive and fast-growing collection of Pre-Trained Models (PTMs) han2021pre; ait2024hfcommunity, compared to other platforms 10.1145/356993410.1145/3643991.3644907. In 2024 alone, HF saw an average of 2,199 new PTMs published per day and over six million daily downloads, based on our analysis of data retrieved through the platform’s public API. Moreover, it is estimated that a new repository is created every 15 seconds  [^1] https://huggingface.co/posts/clem/238420842235482, reflecting trends observed in prior studies that reported rapid increases in repository creation and model contributions 10.1145/3643991.3644898. Despite this growth, the current categorization of HF resources remains largely general-purpose, often targeting ML domains as broad as computer vision or natural language processing. This generic taxonomy fails to address the specific and diverse needs of the Software Engineering (SE) community, whose requirements differ significantly from those of other fields barenkamp2020applications. Importantly, SE goes beyond mere coding, encompassing a set of activities structured around the well-known Software Development Life Cycle (SDLC) ruparelia2010software; sommerville2015. This cycle comprises five main activities: requirements engineering, software design, software implementation, software quality assurance, and software maintenance, thereby covering the entire process from initial specification to long-term system evolution.
@@
For this reason, SE researchers giray2021software and practitioners tan2024challenges; zhao2024empirical face challenges in identifying HF’s ML assets that suit their tasks, while meeting project-specific constraints (e.g., licensing requirements or minimum performance scores). Common issues include missing attributes, discrepancies between reported and actual performance, and risks related to privacy or unethical model behaviour jiang2023empirical. The absence of an SE-specific categorization in HF 10.1145/3661167.3661215 sets a significant barrier, by limiting the efficient application of ML in SE. This gap has severe consequences for SE practitioners and researchers. In the absence of tailored filtering mechanisms, users must manually sift through thousands of PTMs to sample those relevant for their tasks, an effort that is both time-consuming and error-prone, ultimately slowing progress and reducing reproducibility 8804457. Although search engines can facilitate retrieval to some extent, they rely on heterogeneous metadata and lack an SE-specific focus, which favours popular PTMs over those better suited to SE tasks. Moreover, the lack of a task-aligned taxonomy prevents the SE community from systematically identifying relevant resources, evaluating progress, and enabling full automation in the multi-agent environment.
Given the dynamic and evolving nature of both ML and SE 10752650, there is a clear need for a flexible and domain-specific cataloguing system. Such a system would support the effective sampling, selection, and integration of PTMs tailored to the full SDLC. The ability to identify suitable alternatives remains essential for sustainable and adaptable development practices. Moreover, open-source registries such as HF are founded on the principle of choice, offering visibility and reuse opportunities beyond the most popular PTMs. A structured taxonomy can therefore enhance discoverability, guiding users towards PTMs that better fit specific SE needs.
@@
Prior studies have taken important steps toward organizing ML resources for SE, but each presents limitations that leave a significant gap unaddressed. 10.1145/3695988 proposed a taxonomy by categorizing LLM-related papers into SE tasks and SDLC activities through a systematic literature review. However, their approach is grounded in static academic sources and lacks connection to actual open-source registries like HF. 10.1145/3661167.3661215 extracted SE-relevant information from HF model cards using a semi-automated method. Yet, their validation was limited to only three PTMs, and their taxonomy does not align with the SDLC, reducing its applicability. Similarly, yang2024ecosystemlargelanguagemodels curated code-focused LLMs from HF, providing useful insights into that specific subset, but their scope excludes broader SE tasks and lacks an SDLC perspective. citewang2025software surveyed benchmarks for code LLMs and agents around the SDLC. However, their work does not address the classification of other types of PTMs.
@@
In this work, we address these limitations by introducing an automated cataloguing framework for SE tasks across the SDLC. Our approach systematically maps PTMs from HF to a taxonomy of 147 SE tasks, built on prior research and established guidelines. We analyze multiple sources of descriptive content, including model card descriptions, metadata, and abstracts from associated arXiv papers. The pipeline begins with the extraction of candidate task tokens from textual descriptions and applies several processing steps to improve precision. These include outlier detection to remove inconsistent entries, similarity-based deduplication to eliminate near-identical variants, and LLM support to validate that the assigned activity reflects the model’s described purpose. Our main contributions are as follows:
@@
    An automated cataloguing framework that retrieves, processes, and validates PTMs from the HF API, classifying each according to the full SDLC.
@@
    An analysis of SE PTMs, covering their distribution across SE activities and tasks, cross-activity reuse, and their key attributes, including temporal evolution, base models, benchmarks, training datasets, and licenses.
@@
    A validated, publicly available dataset of 2,205 SE PTMs as of March 2025.
@@
Our results offer several key insights. By providing an SDLC-aligned catalogue of PTMs, our work supports researchers and practitioners by automating the identification and selection of the most appropriate PTMs for specific tasks. This automation reduces manual effort, improves consistency in benchmarking, and accelerates the adoption of ML-driven tools within SE workflows, ultimately fostering more effective and targeted automation across the SDLC. In addition, our analysis uncovered that the number of SE PTMs in HF has grown significantly in the past few years, being heavily concentrated on the mid to late stages of the SDLC.
@@
This article is structured as follows. Section 2 reviews related work on PTMs and their application in SE. Section 3 describes the methodology used to catalogue PTMs hosted on HF. Section 4 presents the main results and analysis. Section 5 discusses the findings and explores implications. Section 6 identifies threats to validity; and Section 7 concludes the article and suggests future research directions.
2 Related Work
HF offers a rich ecosystem for conducting empirical studies. Recent repository mining studies have explored the HF ecosystem from several angles to assess its utility and characteristics. ait2025suitability evaluated the suitability and potential of the platform for supporting empirical studies, and jiang2023empirical addressed PTM reuse. Researchers have also investigated standardization and release practices within the ecosystem, including the study of PTM naming conventions by jiang2025see, and the analysis of release transparency, versioning, and naming by ajibode2025towards. Other works have focused on the lifecycle and management of these assets, such as castano2024machine, who conducted a large-scale longitudinal study on how models change over time. While these studies characterize the broader HF landscape, our work specifically targets the identification and cataloguing of SE PTMs.
@@ 2 Related Work
Key aspects of the related work are summarized in Table 1. A systematic literature review conducted by 10.1145/3695988 analyzed 395 research papers from January 2017 to January 2024 and categorized Large Language Models (LLMs) into SE tasks. These tasks were grouped into SE activities according to the SDLC. The study identified 88 distinct SE tasks, revealing that LLMs are most commonly applied in the software development activity, with code generation and program repair being the most prevalent tasks for employing LLMs in software development and maintenance activities, respectively. In contrast, the earlier SDLC phases, such as requirements engineering and software design, are the least explored in the context of LLM applications.
@@
Recent work by 10.1145/3661167.3661215 highlighted the lack of an SE classification of PTMs in HF, as the existing one is specific for ML. To address this gap, they defined six macro-tasks and proposed a semi-automated approach to extract SE-relevant information from model cards. However, their validation was limited to just three PTMs (BERT, RoBERTa, and T5), and their categorization did not align with the SDLC.
@@
Similarly, yang2024ecosystemlargelanguagemodels analyzed the ecosystem of LLMs as of August 2023, curating 366 PTMs from HF for SE. The study also explores the use of LLMs to assist in constructing and analyzing the ecosystem, which increased the model size by 16.5. They focus on code-based LLMs, which represent a more specific scope compared to LLMs for broader SE tasks.
@@
wang2025software examined benchmarks for code LLMs and agents, revealing an imbalance in their distribution across the SDLC. Around 60% of existing benchmarks focus on the software implementation phase, whereas requirements engineering and software design phases are underrepresented.
Complementary to these efforts, ferino2025novice conducted a systematic literature review of 80 studies to investigate LLM adoption among novice developers. Their findings indicated that LLMs are being used across SDLC phases (except for software management), with most of the addressed tasks belonging to software development and little attention on software design. In their study, they organized the activities according to the SDLC, but grouped some of them together (i.e., requirements engineering and software design, software development and software quality assurance, software maintenance).
niu2022deep provided a survey of PTM for code. They categorized 20 CodePTMs across 18 SE tasks and four dimensions: architecture, modality, training tasks, and programming languages. Their taxonomy emphasized the distinction between understanding and generation tasks, highlighting the expanding role of LLMs in automating software implementation activities.
Table 1: Summary of relevant aspects in related work on cataloguing artificial intelligence models for SE
Study Approach githubEmpiricalStandardsdocsstandardsMaster SDLC Classification Information Used Initial Resources Identified Resources 10.1145/3695988 Systematic Review 88 SE tasks across 6 SE activities Paper content 218,765 papers 395 papers 10.1145/3661167.3661215 Repository Mining 6 macro-tasks HF model card 381,240 PTMs Tested on: BERT, RoBERTa, T5 yang2024ecosystemlargelanguagemodels Repository Mining 9 categories HF model documentation 1,566 PTMs 366 PTMs wang2025software Systematic Review 30 SE tasks across 6 SE activities Academic publications 1,347 papers 181 benchmarks from 461 relevant papers ferino2025novice Systematic Review 35 SE tasks across 5 SE activities Paper content 501 studies 80 studies niu2022deep Survey 18 tasks across understanding and generation Model technical details 20 PTMs 20 PTMs This study Repository Mining 58 SE tasks (from a taxonomy of 147 SE tasks) across 5 SE activities HF model card description, metadata, arXiv abstract 1,533,973 PTMs 2,205 PTMs
Beyond academic studies, community platforms such as paperswithcode also offer structured connections between papers and code implementations, offering valuable categorizations of ML tasks  [^2] In July 2025, Papers With Code was integrated into HF as Trending Papers (https://huggingface.co/papers/trending). Although the original website is no longer available, the historical PWC data remains accessible via HF datasets (e.g., https://huggingface.co/datasets/pwc-archive/papers-with-abstracts).. However, these are primarily oriented toward traditional ML domains. Even though the platform includes an SE-related category (i.e., computer code), it combines general ML tasks (e.g., reinforcement learning, semantic segmentation) with SE-specific tasks (e.g., code generation, program repair), limiting its utility for SE-focused research and practice.
@@
Despite the notable advances mentioned above, a comprehensive SE-centric classification framework of PTMs in HF remains absent. Existing efforts, such as those by 10.1145/3695988 and wang2025software, are grounded in static academic literature and are not designed to analyze or update classifications based on the contents of repositories. Other repository-based studies 10.1145/3661167.3661215; yang2024ecosystemlargelanguagemodels do not align their categorization with the SDLC or focus narrowly on a subset of code-centric LLMs. 10.1145/3661167.3661215 take a step toward automation with a semi-automated mapping based on model cards, but their classification remains limited in scope and adaptability. In contrast, this paper adopts an SE perspective to extend the existing taxonomy in 10.1145/3695988, to encompass a broader and more diverse set of PTMs, cataloguing them according to specific SE activities and tasks within the SDLC. By analyzing a substantially larger and broader set of PTMs, this work addresses unique SE-specific requirements not met by previous studies. As a result, we propose a novel and exhaustive classification framework tailored to the SE community.
3 Methodology
In this section, we define the goal and objectives of our study, together with the research questions that guided our investigation. We also provide the rationale behind our approach and describe the different phases of the study.
3.1 Research Objectives
@@
Following the Goal Question Metric (GQM) template caldiera1994goal, our goal is to analyze HF PTMs for the purpose of their automatic classification with respect to their application to SE activities and tasks from the point of view of software engineers in the context of model selection.
@@
This goal is structured around three RQs. The first focuses on key challenges in the automatic cataloguing of PTMs, providing essential information for the subsequent questions:
  • RQ1: What challenges arise when automatically cataloguing SE PTMs in HF?
    - RQ1.1: What is the current documentation status of SE PTMs?
    - RQ1.2: What are the key findings regarding the correct cataloguing of SE PTMs?
@@
Next, we examine the role of PTMs in SE, focusing on their applications across the entire SDLC:
  • RQ2: How are HF PTMs utilized for supporting SE tasks?
@@
    - RQ2.2: How are cross-activity PTMs applied across multiple SE activities?
    - RQ2.3: How are ML tasks related to SE activities?
Lastly, we explore selected characteristics of SE PTMs, including temporal evolution, their underlying base model, associated benchmark scores, training datasets, and licensing information, to provide an in-depth understanding of their properties.
  • RQ3: What key attributes characterize SE PTM in HF?
    - RQ3.1: How has the number of SE PTMs evolved over time?
    - RQ3.2: Which base models have been most frequently utilized for SE over time?
    - RQ3.3: What patterns emerge from the benchmark scores of SE PTMs?
    - RQ3.4: Which training datasets are used to develop SE PTMs?
    - RQ3.5: What types of licenses are applied to SE PTMs?
By addressing these RQs, we aim to understand the current landscape of SE PTMs, thereby characterizing the broader SE PTM ecosystem. Furthermore, by building a database that consolidates these PTMs and their metadata, we seek to support SE researchers and practitioners in discovering suitable PTMs for their projects.
3.2 Research Method
Following the ACM/SIGSOFT Empirical Stardards githubEmpiricalStandardsdocsstandardsMaster for mining repositories, we designed a study to classify PTMs hosted on HF using the pipeline illustrated in Figure 1. The process comprises five main phases. In the first phase (Task Identification), we identified 147 SE tasks from the literature to serve as the basis for the cataloging. In the second phase (Data Collection), we extracted information from each PTM available on the HF platform, including its model card description, metadata, and related arXiv abstracts. The third phase (Data Processing) involved several steps to normalize the collected textual data, detect SE task matches, remove outliers, and identify unique model cards to ensure consistency. In the fourth phase (LLM-Based Validation), we used Gemini 2.0 Flash to refine the catalogue by retaining only PTMs closely aligned to SE. Finally, in the fifth phase (Data Analysis), we analyzed the data gathered throughout the pipeline to answer each of the RQs. The following paragraphs describe each phase in detail.
Figure 1: Pipeline used to extract, catalogue and analyze SE PTMs in HF
This method supports reproducibility, allowing anyone with access to the replication package replication_package to validate the results. Additionally, the pipeline is designed to be easily updated, so researchers can reapply it to classify new PTMs added to the platform using the same method. The inclusion criteria for identifying SE PTMs in HF are summarized in Table 2.
Table 2: Inclusion criteria to identify SE PTMs in HF